OpenStudy (anonymous):

show steps for solving: 434+70y/6-5y=59 Thank you.

OpenStudy (anonymous):

Multiply 59 by (6-5y)

OpenStudy (anonymous):

434+70y=354-295y?

OpenStudy (anonymous):

Right, so now get the y's on one side and the numbers on the other.

OpenStudy (anonymous):

y =-20/91 ?

OpenStudy (anonymous):

Hmm, gone a bit wrong there, I think...

OpenStudy (anonymous):

434+70y= 354-295y minus the 354 on both sides. 80 +70y=-294y minus 70y on both sides.

OpenStudy (anonymous):

-295y not -294 y

OpenStudy (anonymous):

y= -16/73 ?

OpenStudy (anonymous):

\[\huge \checkmark\]

OpenStudy (anonymous):

:) Thank u for helping me.

OpenStudy (anonymous):

ur welcome.

OpenStudy (anonymous):

Watch out for the parenthesis. \[434+70y/6-5y=59\] is different from \[(434+70y/6)-5y=59\] which is different from \[434+(70y/6)-5y=59\] or\[434+(70y/6-5y)=59\] I know it's boring, but it's way better to write all the necessary parenthesis. For the \[(434+70y)/(6-5y)=59\] form, here it goes: \[434+70y=59*(6-5y)\]\[434+70y=(59*6) - (59*5y)\]\[434+70y=354 - 295y\]\[365y = -80\]\[y = -16/73\]\[y \approx -0,219\]

OpenStudy (anonymous):

Thank you. It is very important with the parentheses, and that is a good answer! i'll give you a medal :) :D